Suppose you are given a 45 ∘ −45 ∘ −90 ∘ triangle with hypotenuse of length 2.9 .
Determine the length of the two sides.

Answer by richwmiller(17219)

2x^2=2.9^2
2x^2=8.41
x^2=4.205
x=2.05
The other two sides are 2.05 each